# Compact Cauchy horizons admit constant surface gravity

### v1

@@ Abstract
We prove that in any spacetime dimension and under the dominant energy condition, every totally geodesic connected smooth compact null hypersurface (hence every compact Cauchy horizon) admits a smooth lightlike tangent vector field of constant surface gravity. That is, we solve the open degenerate case by showing that, if there is a complete generator, then there exists a smooth future-directed geodesic lightlike tangent field. The result can be stated as an existence result for a particular cohomological equation. The proof uses elements of ergodic theory, Hodge theory and Riemannian flow theory.
@@ 1 Introduction
Consider a closed, affinely parametrized lightlike geodesic segment $\gamma:[0,\lambda]\to M$, $t\mapsto\gamma(t)$, with $\gamma(\lambda)=\gamma(0)$ and $\dot{\gamma}(\lambda)=a\dot{\gamma}(0)$ for some $a>1$. As observed by Hawking [20, 21], the future-inextendible geodesic $\tilde{\gamma}$ winds around $\gamma$ infinitely many times. On the first loop, it accumulates an affine length $\lambda$; on the second, $\lambda/a$; and so on. The total affine length is therefore $\lambda\sum_{i=0}^{\infty}(1/a)^{i}=\lambda\frac{a}{a-1}$, which is finite. A similar argument in the backward direction shows that the inextendible lightlike geodesic covering $\gamma$ is future incomplete but past complete.
@@
Of course, the general case is far more complex. The generators of the null hypersurface need not close, meaning the hypersurface may exhibit highly intricate dynamics. Nevertheless, as we shall see, under the dominant energy condition and for compact Cauchy horizons, the answer to this question is affirmative.
Understanding the completeness properties of geodesics on compact null hypersurfaces (and their neighborhoods) is of paramount importance. For instance, it could ultimately lead to the removal or relaxation of the null genericity condition in singularity theorems such as those of Hawking and Penrose—a condition that has faced physical criticism [36, Remark 6.22].
The study of compact Cauchy horizons gives also insights into stationary black holes, as their event horizon can be compactified via a standard trick [17]. In fact, the compactification strategy has already served several purposes, from the proof of the regularity of the event horizons of stationary black holes, to the proof of Hawking’s local rigidity theorem [45, 24].
The problem of establishing the constant surface gravity of compact horizons is also motivated by broader considerations. The strong cosmic censorship conjecture (SCC), originally proposed by Penrose [43, 44], asserts that for generic initial data in general relativity, the maximal globally hyperbolic development of a spacetime cannot be extended. This principle encapsulates the idea that, under typical physical conditions, the future evolution of spacetime is uniquely determined by its past. The existence of exact solutions containing Cauchy horizons is understood as a consequence of their non-generic nature. If SCC holds, the formation of compact Cauchy horizons must be highly exceptional, occurring only under very special circumstances.
A precise formulation of this expectation was given by Isenberg and Moncrief [39], who conjectured that any spacetime admitting a non-empty compact Cauchy horizon must exhibit a Killing symmetry—making it non-generic—and that the horizon itself must be a Killing horizon. Their initial proof established this result under three key assumptions: (a) analyticity of the metric and horizon, (b) an (electro-)vacuum spacetime, and (c) closed generators for the horizon.
Subsequent progress has relaxed these restrictions. Notably, it was shown that smoothness of the horizon follows from the smoothness of the metric when the null energy condition holds [31, 35], rendering assumption (a) unnecessary. Under this energy condition, compact Cauchy horizons are smooth, totally geodesic null hypersurfaces, which has become the standard setting for further analysis.
@@
A central question is whether, under the dominant energy condition, $\kappa$ can be normalized to a constant with a suitable choice of $n$, which is a prerequisite for the horizon to be Killing [4, 10]. Constant surface gravity further implies the vanishing of the Lie derivative $L_{n}g=0$ along the horizon, opening the door to extending $n$ to a Killing field via PDE techniques.
@@
Indeed, with the objective of proving constant surface gravity, compact horizons have been classified as degenerate (if at least one generator is complete) or non-degenerate (if at least one generator is incomplete). Isenberg and Moncrief demonstrated a dichotomy under their original assumptions: generators are either all complete (degenerate case) or all incomplete (non-degenerate case) (in which case they are all future-incomplete or all past-incomplete) and ultimately proved that compact horizons are Killing (and, in particular, of constant surface gravity) [39].
@@
Over four decades, their collaborative work [39, 25, 26, 40, 41] progressively weakened assumptions (a) and (c), introducing innovative tools such as Gaussian null coordinates and the ribbon argument. Recent breakthroughs by Bustamante–Reiris and by Gurriaran and the second author [19] established the validity of the dichotomy in the smooth category and under the dominant energy condition. Further, the non-degenerate case was solved in full generality, proving the constancy of $\kappa$ (with $\kappa\neq 0$) under the said assumptions and in any dimensions. Under the additional vacuum assumption, Petersen and Rácz [47, 46, 45] then completed the extension of $n$ to a Killing field, settling the Isenberg–Moncrief conjecture for non-degenerate horizons.
@@
As for the degenerate case, progress in the smooth category has been scarce. However, in a recent work [23] we pointed out the usefulness of Riemannian flow theory and the fact that its implications apply equally well to degenerate and non-degenerate cases. In fact, we were able to obtain a classification result for the possible structures of the horizon. In order to better introduce these findings, we need to be more precise about the definitions.
@@
The achronality property of $H$ will not be important for what follows and so will not be imposed. Anyway, local achronality can be proved by using the existence of convex neighborhoods, and then the spacetime can be reduced to an open neighborhood of $H$ so as to get achronality if desired [19, Prop. 4].
@@
The induced metric on the horizon is denoted $g_{T}$ and termed the transverse metric. Abstracting away from the spacetime, a horizon can be characterized as a triple [^3] For a more general abstract approach to the geometry of null hypersurfaces, see [33]. For the rigidity aspects of null geometry, see [5]. $(H,g_{T},\nabla)$, where:
@@
These conditions imply that the flow of $n$ preserves $g_{T}$ (cf. [30, Thm. 37], [17, Lemma B1], [40], [19, Lemma 7]), in the sense that for $X,Y\in TH$:
@@
of constant dimension, and $L_{X}g_{T}=0$ for all $X\in\Gamma(\ker g_{T})$, then $\ker g_{T}$ is integrable and defines a foliation $\mathcal{F}$ on $H$ [38, Prop. 3.2].
Such a foliation $(H,\mathcal{F})$ is called Riemannian if it arises from a metric $g_{T}$ with the above properties [38, Sec. 3.2]. The metric $g_{T}$ need not be unique. When there is a privileged choice, the pair $(H,g_{T})$ itself might be referred to as a Riemannian foliation.
A Riemannian flow is a Riemannian foliation of oriented 1-dimensional leaves [9]. We can denote with $n$ a non-vanishing, positively oriented vector field tangent to the foliation, unique up to positive scaling. The Riemannian flow condition requires $L_{n}g_{T}|_{TH\times TH}=0$ for some choice of $n$, since the equation $L_{fn}g_{T}|_{TH\times TH}=0$ for any positive function $f$ can be deduced [23]. Not all oriented 1-dimensional foliations admit such a metric [38].
The following proposition connecting the notion of horizon with that of Riemannian flow is not hard to prove [23]:
Proposition 1.2.
@@
We recall that a Riemannian flow is isometric if there exists a Riemannian metric $h$ and a choice of $n$ on $H$ such that $L_{n}h=0$. In fact the pair $(n,h)$ can be chosen so that $h(n,n)=1$ without loss of generality [23, Remark 3]. Not all Riemannian flows are isometric. In [23] we presented several equivalent characterizations of this important property.
The theory of Riemannian foliations began with the work by Reinhart [49] (1959) and systematic treatments have appeared in specialized books, including those by Molino [38] and Tondeur [56]. Reinhart [50] also introduced the basic cohomology. For a Riemannian flow a $k$-form $\alpha$ is basic is $i_{n}\alpha=0$ and $i_{n}{\rm d}\alpha=0$. In other words, basic forms arise as pullback forms of local quotient manifolds (the manifold can be covered by cylinders that admits a quotient under the flow). Note that a 0-th form $f$ is basic if $n(f)=0$, i.e. it is constant over the leaves.
@@
An important result by Carrière establishes that for a Riemannian flow the closure of the orbits are tori [9]. The maximal dimension for a closed tori in the manifold is denoted $k+1$.
By drawing from results by Molino and Carrière [9] we obtained a classification for the structures of a compact horizon in a 4-dimensional spacetime
Theorem 1.4.
@@
      $H$ is diffeomorphic to the lens space [^4] It comprises the case $L(1,0)=S^{3}$ while the case $L(0,1)=S^{1}\times S^{2}$ is included in (ii-b). See [53] for this gluing construction. Note that a $S^{3}$ topology can also appear in case (i) (e.g. Hopf bundle). $L(p,q)$, and the flow is conjugate to that of [23, Example 1].
@@
      $H$ is diffeomorphic to the hyperbolic fibration $T^{3}_{A}$, ($\textrm{tr}A>2$) and the flow is conjugate to one of the flows of [23, Example 2].
@@
This result applies independently of whether the horizon is degenerate or non-degenerate. Under the dominant energy condition non-degenerate horizons are actually isometric [46, 7][23, Cor. 7], and so case (iii-b) is excluded. This also leads to the classification obtained by Bustamante–Reiris for the non-degenerate case via Lie group methods (see also [29] for related results in the analytic case).
@@
Proof.
The statement for dimension 2 is clear as $H$ has dimension 1. In spacetime dimension 3 the horizon has dimension 2 and by Carrière result, the leaves of a Riemannian foliation have closures which are tori and the Riemannian flow restricted to it is conjugate to a linear flow on the torus with dense orbits [8, Thm. V.A.1][9]. Thus either there is a non-closed orbit, which implies that its closure and hence $H$ is a torus with the flow as in (b), or all orbits are closed. In the latter case $H$ is obtained by gluing the two $S^{1}$ sides of a cylinder $S^{1}\times[0,1]$ which can only be done in one way due to the orientation of the generators. This shows that we are in case (a). Both flows preserve the standard flat metric of the torus, hence the flows are isometric. ∎
@@
Theorem 1.6.
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the dominant energy condition. Let $H$ be a compact degenerate horizon, then there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, namely geodesic: $\nabla_{n}n=0$.
Joining with the results of [51, 19]
Corollary 1.7.
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the dominant energy condition. Let $H$ be a compact horizon, then there exists a smooth future-directed lightlike tangent vector field $n$ of constant surface gravity: $\nabla_{n}n=\kappa n$, with $\kappa=cnst$.
The proof is based on elements of Hodge theory and on Riemannian flow theory.
We first provide the proof for the isometric case (Thm. 4.6). In four dimension, there is only one exceptional non-isometric structure which we deal with using some special results. In fact we find two ways of dealing with this special case. Ultimately, then, we prove the theorem for 4-dimensional spacetimes and lower dimensions (Thm. 4.7).
The proof valid for generic dimensions is given later and uses the notion of Carriére neighborhood.
We end this introduction commenting on the problem of the existence of compact degenerate horizons under a vacuum assumption. As noted in [23, Remark 5] they exist and can also be isometric. Isenberg and Moncrief [41] conjectured that they do not exist if additionally the spacetime is analytic and they are Cauchy horizons, $H=H^{+}(S)$. Here, they seem to require that the hypersurface $S$ is spacelike and achronal, which is a somewhat strong form of Cauchy horizon condition with respect to the conventions in [36], according to which, taking as spacetime a neighborhood of $H$ one would always get $H=H^{+}(H)$, see e.g.  [36, Cor. 3.18]. In any case, in our study we shall impose neither analyticity of the metric, nor the Cauchy condition, nor the vacuum assumption, so we do not have reasons to suspect that compact horizons belong to a small family, although results obtained in this work could be used precisely to narrow it down.
@@
As for conventions and terminology, we adopt the notations of [35, 36]. A spacetime $(M,g)$ is a paracompact, time-oriented Lorentzian manifold of dimension $n+1\geq 2$ with signature (,+,,+)(-,+$\dots$,+)( - , + , … , + ). The manifolds $M$ and the metric $g$ are assumed to be smooth. We do not try to work with metrics of regularity $C^{r}$, say $r\geq 3$, because most of the theory of Riemannian flows and Hodge theory on which results we rely has been developed in the smooth category. Still we believe that our main result should hold under much weaker regularity conditions.

### v2

@@ Abstract
We prove that in any spacetime dimension and under the null energy condition, every totally geodesic connected smooth compact null hypersurface (hence every compact Cauchy horizon) admits a smooth lightlike tangent vector field of constant surface gravity. That is, we solve the open degenerate case by showing that, if there is a complete generator, then there exists a smooth future-directed geodesic lightlike tangent field. The result can be stated as an existence result for a particular cohomological equation. The proof uses elements of ergodic theory, Hodge theory and Riemannian flow theory. We emphasize that, remarkably, these results really require only the null energy condition, whereas previous works assumed, already in the Killing or the non-degenerate case, the stronger dominant energy condition.
@@ 1 Introduction
Consider a closed, affinely parametrized lightlike geodesic segment $\gamma:[0,\lambda]\to M$, $t\mapsto\gamma(t)$, with $\gamma(\lambda)=\gamma(0)$ and $\dot{\gamma}(\lambda)=a\dot{\gamma}(0)$ for some $a>1$. As observed by Hawking hawking66b ; hawking73 , the future-inextendible geodesic $\tilde{\gamma}$ winds around $\gamma$ infinitely many times. On the first loop, it accumulates an affine length $\lambda$; on the second, $\lambda/a$; and so on. The total affine length is therefore $\lambda\sum_{i=0}^{\infty}(1/a)^{i}=\lambda\frac{a}{a-1}$, which is finite. A similar argument in the backward direction shows that the inextendible lightlike geodesic covering $\gamma$ is future incomplete but past complete.
@@
Of course, the general case is far more complex. The generators of the null hypersurface need not close, meaning the hypersurface may exhibit highly intricate dynamics. Nevertheless, as we shall see, under the suitable energy conditions and for compact Cauchy horizons, the answer to this question is affirmative.
Understanding the completeness properties of geodesics on compact null hypersurfaces (and their neighborhoods) is of paramount importance. For instance, it could ultimately lead to the removal or relaxation of the null genericity condition in singularity theorems such as those of Hawking and Penrose—a condition that has faced physical criticism (minguzzi18b, , Remark 6.22).
The study of compact Cauchy horizons gives also insights into stationary black holes, as their event horizon can be compactified via a standard trick friedrich99 . In fact, the compactification strategy has already served several purposes, from the proof of the regularity of the event horizons of stationary black holes, to the proof of Hawking’s local rigidity theorem petersen19 ; minguzzi24 .
The problem of establishing the constant surface gravity of compact horizons is also motivated by broader considerations. The strong cosmic censorship conjecture (SCC), originally proposed by Penrose penrose69 ; penrose79 , asserts that for generic initial data in general relativity, the maximal globally hyperbolic development of a spacetime cannot be extended. This principle encapsulates the idea that, under typical physical conditions, the future evolution of spacetime is uniquely determined by its past. The existence of exact solutions containing Cauchy horizons is understood as a consequence of their non-generic nature. If SCC holds, the formation of compact Cauchy horizons must be highly exceptional, occurring only under very special circumstances.
A precise formulation of this expectation was given by Isenberg and Moncrief moncrief83 , who conjectured that any spacetime admitting a non-empty compact Cauchy horizon must exhibit a Killing symmetry—making it non-generic—and that the horizon itself must be a Killing horizon. Their initial proof established this result under three key assumptions:
  • (a)
    the metric and horizon are analytic,
  • (b)
    Einstein’s (electro-)vacuum equations are satisfied,
  • (c)
    the horizon generators are closed.
Subsequent progress has relaxed these restrictions. Notably, it was shown that smoothness of the horizon follows from the smoothness of the metric when the null energy condition holds larsson14 ; minguzzi14d , rendering assumption (a) unnecessary. Under this energy condition, compact Cauchy horizons are smooth, totally geodesic null hypersurfaces, which has become the standard setting for further analysis.
@@
A central question is whether, under the dominant energy condition, $\kappa$ can be normalized to a constant with a suitable choice of $n$, which is a prerequisite for the horizon to be Killing bardeen73 ; chrusciel20 . Establishing constant surface gravity might thus open the way to extending $n$ to a Killing field via PDE techniques.
@@
Indeed, with the objective of proving constant surface gravity, compact horizons have been classified as degenerate (if at least one generator is complete) or non-degenerate (if at least one generator is incomplete). Isenberg and Moncrief demonstrated a dichotomy under their original assumptions: generators are either all complete (degenerate case) or all incomplete (non-degenerate case) (in which case they are all future-incomplete or all past-incomplete) and ultimately proved that compact horizons are Killing (and, in particular, of constant surface gravity) moncrief83 .
@@
Over four decades, their collaborative work moncrief83 ; isenberg85 ; isenberg92 ; moncrief08 ; moncrief20 progressively weakened assumptions (a) and (c), introducing innovative tools such as Gaussian null coordinates and the ribbon argument. Recent breakthroughs by Bustamante–Reiris and by Gurriaran and the second author minguzzi21 established the validity of the dichotomy in the smooth category and under the dominant energy condition. Further, the non-degenerate case was solved in full generality, proving the constancy of $\kappa$ (with $\kappa\neq 0$) under the said assumptions and in any dimensions. Under the additional vacuum assumption, Petersen and Rácz petersen18 ; petersen18b ; petersen19 then completed the extension of $n$ to a Killing field, settling the Isenberg–Moncrief conjecture for non-degenerate horizons.
We can already improve the previous results by observing that constancy of surface gravity in the non-degenerate case (and, as we shall see, also in the degenerate case), holds already under the null energy condition.
The identity for totally geodesic null hypersurfaces (minguzzi21, , Lemma 3)
$Ric(n,X)={\rm d}\omega(n,X),\quad X\in TH,$ (4)
will be useful. Using the dominant energy condition it is possible to derive the condition $Ric(n,\cdot)|_{TH}=0$ which, due to (4), is equivalent to
$i_{n}{\rm d}\omega=0,$
see (minguzzi21, , Lemma 4-5). The last equation is really all that is needed for the application of the ribbon argument, and hence for the dichotomy and the constancy of surface gravity in the non-degenerate case to hold.
However, we have the following simple algebraic fact (we recall that a lightlike vector is a null non-zero vector)
Proposition 1
Let $(M,g)$ be a spacetime and let $n\in T_{p}M$ be a lightlike vector at $p\in M$ such that $R(n,n)=0$ where $R:T_{p}M\times T_{p}M\to\mathbb{R}$ is a symmetric bilinear form at $p$. If $R(v,v)\geq 0$ for every null vector $v\in T_{p}M$ then $R(n,\cdot)=bg(n,\cdot)$ for some constant $b\in\mathbb{R}$.
Proof
Let $N$ be a lightlike vector such that $g(N,n)=-1$, and let $u$ be an arbitrary vector such that $g(n,u)=g(N,u)=0$, hence spacelike. The vector $v=n+xu+x^{2}\frac{1}{2}g(u,u)N$ is lightlike and $R(v,v)=2xR(n,u)+\ldots x^{2}+\ldots x^{3}+\ldots x^{4}$ is a polynomial of fourth-degree whose dominant term for small $x$ is linear. Since $x$ can be chosen arbitrarily small and of any desired sign, the condition $R(v,v)\geq 0$ implies $R(n,u)=0$. Since we already have $R(n,n)=0$ we conclude that for every $w\in\ker g(n,\cdot)$ we have $R(n,w)=0$. Since $\ker R(n,\cdot)\subset\ker g(n,\cdot)$ there is $b\in\mathbb{R}$ such that $R(n,\cdot)=bg(n,\cdot)$ (in fact $b=R(n,N)/g(n,N)$).
Corollary 1
Assume the null energy condition. Over any smooth totally geodesic null hypersurface $H$, if $n\in TH$ is lightlike, we have $R(n,\cdot)|_{TH}=0$ and hence $i_{n}{\rm d}\omega=0$.
Proof
Letting $X=n$ in Eq. (4) we get $R(n,n)=0$, thus by the corollary $R(n,\cdot)|_{TH}=0$, which, again due to Eq. (4), is equivalent to $i_{n}{\rm d}\omega=0$.
Theorem 1.1
All the results in minguzzi21 that use the dominant energy condition actually hold under the weaker null energy condition.
This result is surprising because previous derivations of the constancy of surface gravity for Killing horizons assumed the dominant energy condition bardeen73 ; chrusciel20 . Of course, in this Killing context, the requirement can be weakened to the null energy condition for the reasons stated above. Although the derivation of surface gravity constancy in the Killing case is well-established, this potential improvement has apparently gone unnoticed.
@@
Let us come to the degenerate case. So far progress in the smooth category has been scarce. However, in a recent work minguzzi24c we pointed out the usefulness of Riemannian flow theory and the fact that its implications apply equally well to degenerate and non-degenerate cases. In fact, we were able to obtain a classification result for the possible structures of the horizon. In order to better introduce these findings, we need to be more precise about the definitions.
@@
The achronality property of $H$ will not be important for what follows and so will not be imposed. Anyway, local achronality can be proved by using the existence of convex neighborhoods, and then the spacetime can be reduced to an open neighborhood of $H$ so as to get achronality if desired (minguzzi21, , Prop. 4).
@@
The induced metric on the horizon is denoted $g_{T}$ and termed the transverse metric. Abstracting away from the spacetime, a horizon can be characterized as a triple [^1] For a more general abstract approach to the geometry of null hypersurfaces, see manzano24 . For the rigidity aspects of null geometry, see bekkara06 . $(H,g_{T},\nabla)$, where:
@@
These conditions imply that the flow of $n$ preserves $g_{T}$ (cf. (kupeli87, , Thm. 37), (friedrich99, , Lemma B1), moncrief08 , (minguzzi21, , Lemma 7)), in the sense that for $X,Y\in TH$:
@@
of constant dimension, and $L_{X}g_{T}=0$ for all $X\in\Gamma(\ker g_{T})$, then $\ker g_{T}$ is integrable and defines a foliation $\mathcal{F}$ on $H$ (molino88, , Prop. 3.2).
Such a foliation $(H,\mathcal{F})$ is called Riemannian if it arises from a metric $g_{T}$ with the above properties (molino88, , Sec. 3.2). The metric $g_{T}$ need not be unique. When there is a privileged choice, the pair $(H,g_{T})$ itself might be referred to as a Riemannian foliation.
A Riemannian flow is a Riemannian foliation of oriented 1-dimensional leaves carriere84 . We can denote with $n$ a non-vanishing, positively oriented vector field tangent to the foliation, unique up to positive scaling. The Riemannian flow condition requires $L_{n}g_{T}|_{TH\times TH}=0$ for some choice of $n$, since the equation $L_{fn}g_{T}|_{TH\times TH}=0$ for any positive function $f$ can be deduced minguzzi24c . Not all oriented 1-dimensional foliations admit such a metric molino88 .
The following proposition connecting the notion of horizon with that of Riemannian flow is not hard to prove minguzzi24c :
Proposition 2
@@
We recall that a Riemannian flow is isometric if there exists a Riemannian metric $h$ and a choice of $n$ on $H$ such that $L_{n}h=0$. In fact the pair $(n,h)$ can be chosen so that $h(n,n)=1$ without loss of generality (minguzzi24c, , Remark 3). Not all Riemannian flows are isometric. In minguzzi24c we presented several equivalent characterizations of this important property.
The theory of Riemannian foliations began with the work by Reinhart reinhart59 (1959) and systematic treatments have appeared in specialized books, including those by Molino molino88 and Tondeur tondeur97 . Reinhart reinhart59b also introduced the basic cohomology. For a Riemannian flow a $k$-form $\alpha$ is basic is $i_{n}\alpha=0$ and $i_{n}{\rm d}\alpha=0$. In other words, basic forms arise as pullback forms of local quotient manifolds (the manifold can be covered by cylinders that admits a quotient under the flow). Note that a 0-th form $f$ is basic if $n(f)=0$, i.e. it is constant over the leaves.
@@
An important result by Carrière establishes that for a Riemannian flow the closure of the orbits are tori carriere84 . The maximal dimension for a closed tori in the manifold is denoted $k+1$.
By drawing from results by Molino and Carrière carriere84 we obtained a classification for the structures of a compact horizon in a 4-dimensional spacetime
Theorem 1.2
@@
      $H$ is diffeomorphic to the lens space [^2] It comprises the case $L(1,0)=S^{3}$ while the case $L(0,1)=S^{1}\times S^{2}$ is included in (ii-b). See saveliev12 for this gluing construction. Note that a $S^{3}$ topology can also appear in case (i) (e.g. Hopf bundle). $L(p,q)$, and the flow is conjugate to that of (minguzzi24c, , Example 1).
@@
      $H$ is diffeomorphic to the hyperbolic fibration $T^{3}_{A}$, ($\textrm{tr}A>2$) and the flow is conjugate to one of the flows of (minguzzi24c, , Example 2).
@@
This result applies independently of whether the horizon is degenerate or non-degenerate. Under the dominant (actually null, as we have shown previouslyin Thm. 1.1) energy condition non-degenerate horizons are isometric petersen18b ; bustamante21 (minguzzi24c, , Cor. 7), and so case (iii-b) is excluded. This also leads to the classification obtained by Bustamante–Reiris for the non-degenerate case via Lie group methods (see also kroenke24 for related results in the analytic case).
@@
Proof
The statement for dimension 2 is clear as $H$ has dimension 1. In spacetime dimension 3 the horizon has dimension 2 and by Carrière result, the leaves of a Riemannian foliation have closures which are tori and the Riemannian flow restricted to it is conjugate to a linear flow on the torus with dense orbits (carriere81, , Thm. V.A.1)carriere84 . Thus either there is a non-closed orbit, which implies that its closure and hence $H$ is a torus with the flow as in (b), or all orbits are closed. In the latter case $H$ is obtained by gluing the two $S^{1}$ sides of a cylinder $S^{1}\times[0,1]$ which can only be done in one way due to the orientation of the generators. This shows that we are in case (a). Both flows preserve the standard flat metric of the torus, hence the flows are isometric.
@@
Theorem 1.4
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the null energy condition. Let $H$ be a compact degenerate horizon, then there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, namely geodesic: $\nabla_{n}n=0$.
Joining with the results of reiris21 ; minguzzi21
Corollary 2
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the null energy condition. Let $H$ be a compact horizon, then there exists a smooth future-directed lightlike tangent vector field $n$ of constant surface gravity: $\nabla_{n}n=\kappa n$, with $\kappa=cnst$.
The proof is based on elements of ergodic theory, Hodge theory and Riemannian flow theory.
We first provide the proof for the isometric case (Thm. 4.4). In four dimension, there is only one exceptional non-isometric structure which we deal with using some special results. In fact we find two ways of dealing with this special case. Ultimately, then, we prove the theorem for 4-dimensional spacetimes and lower dimensions (Thm. 4.5).
The proof valid for generic dimensions is given later and uses the notion of Carrière neighborhood.
We end this introduction commenting on the problem of the existence of compact degenerate horizons under a vacuum assumption. As noted in (minguzzi24c, , Remark 5) they exist and can also be isometric. Isenberg and Moncrief moncrief20 conjectured that they do not exist if additionally the spacetime is analytic and they are Cauchy horizons, $H=H^{+}(S)$. Here, they seem to require that the hypersurface $S$ is spacelike and achronal, which is a somewhat strong form of Cauchy horizon condition with respect to the conventions in minguzzi18b , according to which, taking as spacetime a neighborhood of $H$ one would always get $H=H^{+}(H)$, see e.g.  (minguzzi18b, , Cor. 3.18). In any case, in our study we shall impose neither analyticity of the metric, nor the Cauchy condition, nor the vacuum assumption, so we do not have reasons to suspect that compact horizons belong to a small family, although results obtained in this work could be used precisely to narrow it down.
@@
As for conventions and terminology, we adopt the notations of minguzzi14d ; minguzzi18b . A spacetime $(M,g)$ is a paracompact, time-oriented Lorentzian manifold of dimension $n+1\geq 2$ with signature (,+,,+)(-,+$\dots$,+)( - , + , … , + ). The manifolds $M$ and the metric $g$ are assumed to be smooth. We do not try to work with metrics of regularity $C^{r}$, say $r\geq 3$, because most of the theory of Riemannian flows and Hodge theory on which results we rely has been developed in the smooth category. Still we believe that our main result should hold under much weaker regularity conditions.